A wave has a frequency of 475 Hz. What is its period?
Answer #1

period = 1/frequency

= 1/475

=2.1* 10^-3 s ans
Answer #2

period = 1/frequency
   =1/475
   =0.0021 s
   =2.1 ms
Answer #3

T=1/f=1/475=2.105*10^-3 s=2.103 ms
Answer #4

T = 1/f

= 1/475

= 2.1*10^-3 s or 2.1 ms